OpenStudy (anonymous):

The length of a rectangle is 2 units and its width is square root of 2 unit. Is the area of the rectangle rational or irrational? Justify your answer.

OpenStudy (anonymous):

i could use the help please ill award a medal

OpenStudy (anonymous):

just logged on let me work it out

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Irrational, and I assume you want the proof of why its irrational?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

im guessing because the area is not a whole number

OpenStudy (anonymous):

\[\sqrt{2} \] is irrational. Proof by way of contradiction. assume that \[\sqrt{2}\] is rational. Then is can be expressed as \[\frac{ a }{ b }\] where \[a,b \in \integer\] , \[b \neq 0\] when a and b have no common factors.

OpenStudy (anonymous):

is the first part let me finish

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

\[\sqrt{2}=\frac{ a }{ b }\] \[2= [\frac{a}{b}]^2\] \[2b^2=a^2\] Because a is an integer it can be expressed as a=2k where k is some integer. by substituting 2k in the equation for a we have \[2b^2=(2k)^2\] \[2b^2=4k^2\] but this implies that a and b share a common factor of 2. which violates the premise that a and b share no common factors. therefore \[\sqrt{2}\] must be irrational. this leads us to conclude that \[2\sqrt{2}\] must also be irrational as any rational number times an irrational number is irrational Q.E.D.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

do you think you can help with another question

OpenStudy (anonymous):

I can try.

OpenStudy (anonymous):

okay ill show you

OpenStudy (anonymous):

A scientist is studying the growth of a particular species of plant. He writes the following equation to show the height of the plant f(n), in cm, after n days. f(n) = 10(1.02)n Part A: When the scientist concluded his study, the height of the plant was approximately 11.04 cm. What is a reasonable domain to plot the growth function? (4 points) Part B: What does the y-intercept of the graph of the function f(n) represent? (2 points) Part C: What is the average rate of change of the function f(n) from n = 1 to n = 5 and what does it represent? (4 points)

OpenStudy (anonymous):

sorry i have big problems with functions i would really appreciate if you could help

OpenStudy (anonymous):

Well i can try it, but what math class is this for, itll help me narrow down the approach your gonna want to use?

OpenStudy (anonymous):

working with functions

OpenStudy (anonymous):

oh im taking algebra

OpenStudy (anonymous):

perfect, and one question, is the number (1.02) supposed to be a power by chance or are all of the given numbers 10, 1.02, n simply numbers?

OpenStudy (anonymous):

umm its just the number that represents how much the plant grows and 10 represents the height of plant

OpenStudy (anonymous):

f(n) = 10(1.02)n

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

a reasonable range would be \[0 \le y \le 2\] because the equation \[y=10(1.02)n\] can be rewritten as \[y= 10.2n\] given the final value of 11.04 we substitute that into the equation for y \[11.04=10.2n\] \[\frac{11.04}{10.2} =n\] 1.08235.. = n so the researcher quit after one day. the domain is the number of days, so on a graph the number of days would be 0 to 2

OpenStudy (anonymous):

okay so thats for part a

OpenStudy (anonymous):

the y intercept repressents the height of the plant at the start of the experiment. which is part b

OpenStudy (anonymous):

oh okay and for part c it would be

OpenStudy (anonymous):

doing the math now, give me a second

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

The rate of change is the slope of the line from the points (1,10.2) to (5,51) [i got the y coordinate by plugging in the n values 1 and 5 into the equation and solving for y. the slope is the rise over run of the points so: \[\frac{51-10.2}{5-1}\] \[\frac{40.8}{4}\] \[10.2\] this number represents the average amount of growth per day of the plant

OpenStudy (anonymous):

thank you so much i have questions but i dont bother you since you already helped me alot thanks for your time